The average mark of candidates in an aptitude test was 128.5 with a standard deviation of 8.2. Three scores extracted from the test are; 148, 102, 152. What is the average of the extracted scores that are extreme values (outliers)?
1
Expert's answer
2020-06-02T18:26:48-0400

"\text{According to Z- score }\\
\text{any data value with a z-score greater }\\
\text{ than -2 or less than 2 as an outlier}\\
Z=\frac{x- \mu}{\sigma}\\
Z_{1}=\frac{148- 128.5}{8.2}=2.38\\
Z_{2}=\frac{102- 128.5}{8.2}=-3.23\\
Z_{3}=\frac{152- 128.5}{8.2}=2.86\\
\text{ so all of them are outliers.}\\
\text{the average of the extracted scores that }\\
\text{are extreme values (outliers)}
\\=\frac{148+102+152}{3}=134"